A block of mass 3.0 π‘˜π‘” slides with uniform velocity down a plane inclined 25o with the horizontal. If the angle of inclination is increased to 40o , what will be the acceleration of the block?Β 


1
Expert's answer
2021-09-30T14:24:17-0400

Given:

"\theta_1=25^{\circ}"

"\theta_2=40^{\circ}"


The friction cooficient

"\mu=\tan\theta_1=\tan 25^{\circ}=0.466"

The acceleration of the object on the inclined plane

"a=g(\sin\theta_2-\mu\cos\theta_2)"

Hence,

"a=9.81*(\sin40^{\circ}-0.466*\cos40^{\circ})=2.80\:\rm m/s^2"